OpenStudy (anonymous):

steps to solving : 3x-4y-4z=8 4x+2y-2z=11 -5x+8y+3y=-9

OpenStudy (helder_edwin):

1st eq + 2(2nd eq) 2(1st eq) + 3rd eq

OpenStudy (anonymous):

ok i thought it was 4(1st eq) + -3(2nd eq) then, 5(2nd eq) + 4(3rd eq)

OpenStudy (helder_edwin):

that would also work

OpenStudy (helder_edwin):

what i posted would eliminate y what u posted would eliminate x

OpenStudy (anonymous):

i try it ddnt get the right answer i checked my work and i dont see any mistakes and yes i was tryin to solve for y and z first and then find x . I got y = 0

OpenStudy (helder_edwin):

which way do u wanna go? yours or mine?

OpenStudy (anonymous):

I'm tryin yours right now to see what i get.

OpenStudy (helder_edwin):

ok let's see 1st eq + 2*(2nd eq) 3x-4y-4z=8 8x+4y-4z=22 ------------- 11x-8z=30 2*(1st eq) + 3rd eq 6x-8y-8z=16 -5x+8y+3y=-9 --------------- x-5z=7

OpenStudy (anonymous):

yes thats what i got so far

OpenStudy (helder_edwin):

so x=7+5z then 11(7+5z)-8z=30 77+55z-8z=30 47z=-47 z=-1 x=7+5(-1)=2 3(2)-4y-4(-1)=8 6-4y+4=8 -4y=-2 y=1/2

OpenStudy (helder_edwin):

did u get the same?

OpenStudy (anonymous):

yes i did and i got x=2 after pluging y and z in is that correct?

OpenStudy (helder_edwin):

look at what i did. u should get first either x and z, or z and x, and then y

OpenStudy (anonymous):

Yes, i solved for x and z and then solved for y , i know understand how the process in this goes, thanks.

OpenStudy (helder_edwin):

u r welcome